OpenStudy (anonymous):

PLEASE HELP ANYONE A prism of height 12" has a rhombus with diagonals 6" and 8" for a base. Find the volume and the lateral area. V = cu. in. L.A. = sq. in.

OpenStudy (anonymous):

Volume=300 Dunno what la means.

OpenStudy (anonymous):

Draw a picture, with your diagonals running accrost the front. See how they intersect? On each side of the intersection is half the diagonal length, and the lines form right angles

OpenStudy (anonymous):

sending the answer....wait....

OpenStudy (anonymous):

So u have 4 different right triangles, with lengths of 3,4, then obviously 5.

OpenStudy (anonymous):

so do base times width (5*5) times height(12) to find volume

OpenStudy (anonymous):

volume of prism = area of base * height here base is a rhombus with length of diagonals given area of rhombus = half the product of lengths of diagonals 6 in * 8 in 48 sqin so, area of base = half the product of diagonals = --------- = -------- = 24 sqin 2 2 Volume = 24 sqin * 12 in = 288 cubic inch

OpenStudy (anonymous):

how do i do La ?

OpenStudy (anonymous):

wait pls...

OpenStudy (anonymous):

for this we first need to find the length of the side of the base rhombus...

OpenStudy (anonymous):

diagonals of a rhombus are perpendicular bisectors of each other. so, when the two diagonals bisect each other, they form a right angled triangle whose hypotenuse is the side of the rhombus and the perpendicular and base are half of each of the diagonals..... so (side of rhombus)² = 3² + 4² = 9 + 16 = 25 so side or rhombus = √25 = 5 so each side of rhombus is 5 inch rhombus has four sides and so on each side we will have a vertical rectangular face with dimensions 5inch * 12 inch so LA = 4* [5in * 12in] = 4 * 60 sqin = 240 sqin